AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 00:42:54 2025
Author
*
:
Subject
*
:
01.07 DAQ continues file G22-88/R1_46 background alpha DSSSD bias OK - attachment 1 FEE64 temps OK - attachment 2 ADC data item stats - attachment 3 aida15 asic #2 u/s - rate c. 266k aida11 1*L spectra - attachment 4 ADC data items/s Merger, Tape Server & Merger stats - attachments 5-7 data rate to disk c. 2Mb/s dominated by aida15 asic 2 10.05 DAQ continues file G22-88/R1_79 background alpha DSSSD bias OK - attachment 8 ambient temperature 16.8 deg C FEE64 temps OK - attachment 9 ADC data item stats - attachment 10 aida15 asic #2 u/s - rate c. 266k aida11 1*L spectra - attachment 11 ADC data items/s Merger, Tape Server & Merger stats - attachments 12-14 data rate to disk c. 2Mb/s dominated by aida15 asic 2 10.12 per FEE64 Stat spectra - attachment 15 missing channels aida03, aida04, aida05 & aida15 Check ASIC load aida03, aida04, aida05 & aida15 13:25 Overnight leakage current plot - att 16 13.38 DAQ continues file G22-88/R1_92 background alpha DSSSD bias OK - attachment 17 ambient temperature 16.8 deg C FEE64 temps OK - attachment 18 ADC data item stats - attachment 19 aida15 asic #2 u/s - rate c. 266k aida11 1*L spectra - attachment 20 ASIC check load at 10.12 should result in 1-2 hits per channel in the Stat spectra aida03, aida04, aida05 and aida15 Channels which continue to show zero counts likely indicate a FEE64-adaptor PCB connection issue Channels which now show 1, or more, counts cf. attachment 15 likely correspond to channels now working following the check ASIC load ADC data items/s Merger, Tape Server & Merger stats - attachments 21-23 data rate to disk c. 2Mb/s dominated by aida15 asic 2 13.48 Stopped DAQ file G22-88/R1_93 DSSSD bias OFF FEE64 power OFF 16.00 Installed FEE64 power cable to aida16 Replaced aida15 with new FEE64 MAC 41:f6:5a Update dhcpd.conf Update MIDAS config to include aida16 FEE64 power ON All system wide checks OK DSSSD bias OK - attachment 24 FEE64 temperatures OK - attachment 25 ADC data item stats - attachment 26 aida15 rate changes from c. 260k to c. 25k ( 2x 'hot' channels, 1x channel missing, all other channels OK ) per FEE64 Rate spectra - attachment 27 per FEE64 1.8.L spectra - attachments 28-29 pulser peak width aida11 44 ch FWHM, aida14 57 ch FWHM per FEE64 1.8.W spectra - 20us FSR - attachments 30-31 Update NewMerger startup script and Options for 16x FEE64s - see attachments 32-33 16.20 file G22-88/R2 Pulser 50Hz 16.28 DAQ stop Pulser OFF All histograms zero'd 16.33 DAQ starts file G22-88/R3 alpha background ADC data item stats OK - attachment 34 NewMerger, TapeServer and Merger stats - attachments 35-38 16.50 Pulser peak width spreadsheet for file R2 - attachment 39 20.32 DAQ continues file G22-88/R3_9 DSSSD bias OK - attachment 40 FEE64 temperatures OK - attachment 41 ADC data item stats - attachment 42 aida15 rate c. 25k ( 2x 'hot' channels, 1x channel missing, all other channels OK ) per FEE64 Stat spectra - attachment 43 Merger, Merger stats and TapServer - attachments 44-46 Disc data items - attachment 47 aida15 disc rate c. 8k Disable aida15 disc ( fast comparator ) outputs for asic#3 - attachment 48 Disc data items- attachment 49 all FEE64s rate << 1Hz Update NewMerger startup script and Options for 16x FEE64s - see attachments 32-33
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown